40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 9 23:17:47 2010, kiwamu, Update, SUS, beam pointing has been done 
    Reply  Fri Dec 10 09:58:57 2010, Aidan, Update, SUS, beam pointing has been done 
       Reply  Sat Dec 11 00:41:55 2010, kiwamu, Update, IOO, alignment of IP_ANG is done 
          Reply  Tue Dec 14 01:04:23 2010, kiwamu, Update, SUS, alignment of ITMs and PRM done 
             Reply  Tue Dec 14 15:15:19 2010, kiwamu, Update, SUS, put ETMY and aligned the beam 
                Reply  Tue Dec 14 22:42:56 2010, kiwamu, Update, SUS, installed the last TT . And some more progresses 
                   Reply  Thu Dec 16 01:23:43 2010, kiwamu, Update, SUS, SRM OSEMs installed DSC_2757_ss.jpg
Message ID: 4040     Entry time: Fri Dec 10 09:58:57 2010     In reply to: 4039     Reply to this: 4044
Author: Aidan 
Type: Update 
Category: SUS 
Subject: beam pointing has been done 

Good news. I feel multi-chromatic-locking success is just around the corner.

By the way, there's a new presentation on the DCC from the ANU group where they've locked a short single cavity with both colors - G1000735:

https://dcc.ligo.org/cgi-bin/private/DocDB/ShowDocument?docid=14040

 

 

Quote:

[Koji, Osamu and Kiwamu]

 We aligned the beam axis pointing down to both X and Y arm.

Now the beams are hitting the centers of both ETMX and ETMY.

Amazingly Osamu made X arm flashing by aligning the cavity.


 

ELOG V3.1.3-